Please show steps for solving for "k".
26=1.30(1+0.12)/k-0.12 = 1.456/k-0.12

Answer by ewatrrr(24785)

 
Hi,
26=1.30(1+0.12)/(k-0.12)
26 = 1.456/(k-0.12)
k -0.12 = 1.456/26
k = 1.456/26 + .12 = .176